A risky portfolio has an expected return of 12% and standard deviation of 25%. Given a risk free rate of 3%, what percentage of a clients portfolio should be allocated to the risky portfolio if the client has a risk aversion of 4?
Answer #1

Given that,

Expected return on risky portfolio E(r) = 12%

Standard deviation SD(r) = 25%

Risk free rate Rf = 3%

risk aversion of client, A = 4

So, weight of risky asset in a optimal complete portfolio is calculated using formula

w = (E(r) - Rf)/(A*SD(r)^2) = (0.12 - 0.03)/(4*0.25*0.25) = 0.36

So client should allocate 36% of his portfolio to risky portfolio.